Ask
your own question, for FREE!
Mathematics
11 Online
Determine the probability that at least one head appears in three flips of a fair coin. Round your answer to one decimal place.
Still Need Help?
Join the QuestionCove community and study together with friends!
is it .875
all 8 options are HHH,HHT,HTH, HTT, THH THT TTH TTT. (8 in total) seven do have atleast one head so 7/8.
or u cud say that P(3 heads) = 1/8 P(at least one head) = 1- p(3 heads) = 1- 1/8 = 7/8
= 0.875 = 0.9 (rounded off to one decimal place)
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
13 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal